# Unifying Bias and Unfairness in Information Retrieval: A Survey of Challenges and Opportunities with Large Language Models

### v1

@@ 3.2.2. Popularity Bias
$\bullet$ Popularity Bias. LLM-based IR models tend to prioritize candidate documents or items with high popularity levels.
Popularity Bias: LLM-based IR models tend to prioritize candidate documents or items with high popularity levels.
@@ 3.2.3. Instruction-Hallucination Bias
Instruction-Hallucination Bias: Content generated by LLM-based IR models may deviate from the instructions provided by users.
@@ 3.2.4. Context-Hallucination Bias
Context-Hallucination Bias: LLMs-based IR models may generate content that is inconsistent with the context.
There have been many studies showing that LLMs run the risk of generating content that is inconsistent with the context, especially in scenarios where the context is very long and multi-turn responses are needed (Liu et al., 2024; Qiu et al., 2024; Bai et al., 2024; Li et al., 2023f). When using LLMs as recommenders in scenarios with complex context such as multiple rounds of conversation history and user portrait information, LLMs may give the items that are contradictory with the conversation history and user preferences. This emerges when LLMs cannot understand the context or fail to maintain consistency throughout the conversation (Zhang et al., 2023d), which is mainly because LLMs still have limitations in processing long texts (Liu et al., 2024). Therefore, the main research to mitigate the context-hallucination bias focuses on improving the memory and processing capabilities of LLMs for long texts (Xu et al., 2023d; Chen et al., 2023c; Wang et al., 2024).
@@ 3.3. Bias in Result Evaluation
Selection Bias: LLM-based evaluators may favor the responses at specific positions or with specific ID tokens.
For example, Zheng et al. (2023c) have demonstrated that gpt-3.5-turbo exhibits a preference for choice “C”, while llama-30B shows a preference for choice “A” across various benchmarks. Other works have revealed a common tendency among LLMs to favor responses positioned at the first position  (Wang et al., 2023b; Pezeshkpour and Hruschka, 2023; Koo et al., 2023; Zheng et al., 2024; Chen et al., 2024). Selection bias may stem from an imbalance in how answers of different positions or with distinct ID options are represented in the training data (Zheng et al., 2024, 2023c). Further investigation reveals that this bias tends to amplify when LLM-based evaluators are uncertain about the prediction between the top-ranked choices(Wang et al., 2023b; Zheng et al., 2023c).
To address selection bias, several approaches have been explored: (1) Prompting: Simple prompt-based methods, such as incorporating few-shot examples or employing chain-of-thought and reference-guided judgment, have been proposed (Zheng et al., 2024; Pezeshkpour and Hruschka, 2023; Kim et al., 2023). However, these methods have shown limited effectiveness. (2) Data Augmentation: Strategies like position or token switching aim to eliminate selection bias by diversifying the evaluation context (Zheng et al., 2024; Chern et al., 2024; Zheng et al., 2023b; Wang et al., 2023b; Pezeshkpour and Hruschka, 2023). While these methods can enhance the robustness of evaluations, they are often time-consuming and costly. (3) Rebalancing: Zheng et al. (2023c) utilized probability decomposition techniques to estimate the prior distribution of specific positions or tokens associated with responses, which helps in aligning evaluations closer to objective standards. Wang et al. (2023b) have developed a calibration framework that integrates Human-in-the-Loop to calculate balanced position diversity entropy for final selection. Despite the availability of these strategies, effectively mitigating selection bias in LLM-based evaluators remains a challenge, requiring further research for more efficient and cost-effective solutions.
@@ 3.3.1. Style Bias
Style Bias: LLM-based evaluators may favor the responses with specific styles (e.g., longer responses).
For instance, several studies have identified a clear preference in LLMs for longer responses over shorter ones, emphasizing form over the actual quality of the content (Wu and Aji, 2023; Zheng et al., 2024; Liu et al., 2023b; Saito et al., 2023). Moreover, Chen et al. (2024) have observed an inclination towards content with visually engaging elements, such as emojis or references, even when such content may include factual errors. Huang et al. (2024) suggest that this bias may stem from the training process of LLMs, which emphasizes generating fluent and verbose responses, thereby inadvertently leading them to prefer these characteristics when employed as evaluators.
@@ 3.3.2. Egocentric Bias
With LLMs being extensively utilized in the development of IR models, egocentric bias has emerged as a new bias during the automated evaluation conducted by LLMs (Liu et al., 2023b, a; Koo et al., 2023; Xu et al., 2024d), which can be defined as follows:
Egocentric Bias: LLM-based evaluators prefer the responses generated by themselves or LLMs from the same family.
@@
Data Collection User Unfairness (Ghanbarzadeh et al., 2023; Zhang et al., 2023a; Xu et al., 2023e; Lu et al., 2020; Dixon et al., 2018; Sun et al., 2022b; Ung et al., 2021) (Raffel et al., 2020; Ngo et al., 2021) (Orgad and Belinkov, 2022; Deldjoo and di Noia, 2024) (Qian et al., 2019; Bordia and Bowman, 2019; Huang et al., 2019) (Fang et al., 2024)
@@
Item Unfairness (Grosse et al., 2023) (Shi et al., 2023a; Akyürek et al., 2022) (Wang et al., 2023h; Zhang et al., 2023b, e; Wang et al., 2023c; Sander et al., 2024)
@@ 4. Cause and Mitigation of Unfairness
4.1. Fairness Concept
Sociological researches acknowledge multiple cultural variations in perceptions of fairness (Tyler and Allan Lind, 2002; Tyler and Smith, 1995). In IR systems, achieving fairness often entails ensuring that the retrieved documents or recommended items align with cultural values (Mehrabi et al., 2021; Li et al., 2023a), including principles such as gender equality (Fang et al., 2024; Zhang et al., 2023a; Xu et al., 2023e; Li et al., 2023a), addressing disadvantages (Xu et al., 2023a; Patro et al., 2020), and avoiding discriminatory language (Wen et al., 2023; Gallegos et al., 2023).
Researchers have revealed that various multi-stakeholders involved in IR systems (Abdollahpouri et al., 2020; Chen et al., 2023a), such as users and items, often have distinct perspectives on fairness considerations. In the realm of IR, user fairness and item fairness are often associated with two sociological concepts: equality and distributive justice (Xu et al., 2023a).
User fairness aligns with the concept of equality (Matsumoto and Juang, 2016), which presupposes that every user is inherently equal, and thus, different users should receive identical support:
User Fairness: Everyone should be treated the same and provided the same resources to succeed. This implies that the IR systems should deliver equitable and non-discriminatory information services to different users.
Conversely, item fairness corresponds to the concept of distributive justice, as proposed by Rawls (1958), which operates under the assumption that distinct items inherently vary, and thus, weaker items should receive more support:
Item Fairness: The resources should be equally distributed based on needs. This implies that the IR systems should afford more opportunities (e.g., exposures) to weaker items, striving to equalize the opportunities across diverse items.
Next, we will show how user unfairness and item unfairness happen in the three stages of the integration of LLMs into IR systems.
@@ 4.2.1. User Unfairness
In the context of user unfairness, one primary cause stems from the existing taxonomic, discriminatory, and offensive content in the training data, disproportionately affecting specific groups. The inclusion of these contents within the existing material can be attributed to historical and cultural reasons (Ntoutsi et al., 2020; Zhuo et al., 2023; Beukeboom and Burgers, 2019), or they may be generated by LLMs (Fang et al., 2024). For example, when training large language models, it is unfortunately common to encounter instances with discriminatory content (Beukeboom and Burgers, 2019; Blodgett et al., 2016; Abid et al., 2021; Fang et al., 2024; Das et al., 2024). The existence of discrimination against certain groups can also stem from unbalanced data collection, where insufficient representation of diverse perspectives leads to unfair outcomes (Ghanbarzadeh et al., 2023; Mehrabi et al., 2021). The presence of unbalanced data can contribute to the perpetuation of historical and cultural stereotypes (Ellemers, 2018; Haines et al., 2016; Hurwitz and Peffley, 1997; Ladhak et al., 2023) or systematic influences (Chen et al., 2023a).
In previous works, various methods have been employed to mitigate unfairness during data collection by redistributing the existing item or document corpus. Generally, to address the imbalance in the original dataset against certain groups, some work (Ghanbarzadeh et al., 2023; Zhang et al., 2023a; Xu et al., 2023e; Lu et al., 2020) create matched pairs (e.g., male or female) to ensure a more equitable dataset and other methods (Dixon et al., 2018; Sun et al., 2022b; Ung et al., 2021) add non-toxic examples for groups. Other approaches (Orgad and Belinkov, 2022; Deldjoo and di Noia, 2024) suggest the use of down-weighting samples containing social group or discriminated information as a re-sampling strategy. Moreover, other studies (Raffel et al., 2020; Ngo et al., 2021) propose to filter out and remove discriminated or taxonomic content from web-scale datasets. Instruction fine-tuning or RLHF has also been shown to be effective in promoting fairness (Touvron et al., 2023; Lu et al., 2020).
@@ 4.2.2. Item Unfairness
For item unfairness, one primary cause is likely unbalanced data collection, where insufficient representation of certain items leads to disparities in the assessment or information retrieval process (Jiang et al., 2024a). Another reason raised is that LLMs cannot only retrieve existing items but also generate new items, contributing to the potential introduction of novel content and perspectives (Das et al., 2024; Laverghetta Jr and Licato, 2023; Guenole et al., 2024; Götz et al., 2023; Zu et al., 2023). However, these newly generated items may still encounter discrimination issues (Laverghetta Jr and Licato, 2023; Guenole et al., 2024).
@@ 4.3.1. User Unfairness
When adapting LLMs as information retrievers, researchers have observed that the extensive knowledge gained during pre-training may introduce risks of user unfairness (Xu et al., 2023e; Zhang et al., 2023a; Ghanbarzadeh et al., 2023; Deldjoo and di Noia, 2024), highlighting the need for careful consideration and mitigation strategies in deploying such models. Studies (Zhang et al., 2023a; Deldjoo and di Noia, 2024; Ghanbarzadeh et al., 2023) have shown that utilizing explicit user-sensitive attributes like gender or race in LLMs may lead to the generation of discriminated recommendation results or unfair answers to specific questions. Moreover, it has been observed that LLMs can learn implicit attributes, such as user names and email addresses, and utilize them to generate discriminated content (Xu et al., 2023e; Wolfe and Caliskan, 2021).
To address user unfairness in the model development phase, prior research has suggested mitigating unfairness during the fine-tuning process as an effective strategy. Some studies (Deldjoo and di Noia, 2024) investigate how different intersectional prompts affect recommendation fairness and UP5 (Hua et al., 2023; Yang et al., 2023b) propose to conduct prompt tuning to get an effective fair-aware prompt. Han et al. (2021); Zayed et al. (2023) proposes to set different weights of loss for different samples containing discriminated content. Meanwhile, various studies (Liu et al., 2019; Woo et al., 2023; Park et al., 2023; Wang et al., 2023a; Garimella et al., 2021; Zhou et al., 2023) explore the incorporation of a fairness-aware regularizer to assist LLMs-based models in generating more equitable content. Wang et al. (2021) propose to remove unfair information from LLMs-based embeddings by generating adversarial examples. There are also some works (Wang et al., 2021; Meade et al., 2023; Schramowski et al., 2022; Shuster et al., 2022) that recommend employing a filtering-list approach or comparing model outputs with safe samples to proactively prevent the inclusion of discriminatory words and enhance the fairness of the generated content. The most common ways are through RLHF (Ouyang et al., 2022) and RLAIF (Bai et al., 2022).
@@ 4.4.2. Item Unfairness
In evaluating item unfairness, when turning from discriminant style to generation style, a primary concern arises from attributing credit to the generated items (Shi et al., 2023a; Grosse et al., 2023; Akyürek et al., 2022; Wang et al., 2023c; Sander et al., 2024), as achieving item fairness necessitates tracing this credit back to the item provider for a comprehensive assessment.
@@
To tackle these challenges, several studies (Wang et al., 2023h; Zhang et al., 2023b, e) have developed IR agents that use world knowledge to fairly distribute item credits, aiming to reduce biases and enhance fairness. Shi et al. (2023a) employs the MIN-KProb technique to check whether an item exists in the training data. Meanwhile, Grosse et al. (2023); Akyürek et al. (2022) leverage influence functions and embedding similarities for item credit assessment. Additionally, other research (Wang et al., 2023c; Sander et al., 2024) explores watermarking techniques for item tracking.
5. Challenges and Future Directions
Despite the considerable attention given to this topic, we identify some important problems for further exploration in this section.
5.1. Biases and Unfairness in IR Feedback Loops
In real IR systems, the interaction between users, models, and information is dynamic, forming feedback loops that evolve and impact each other over time. These loops can significantly shape user perceptions, behaviors, and preferences based on the information they are exposed to. This interaction, in turn, influences the training data for IR models, creating a cycle that may reinforce and amplify existing biases and unfairness. Novel strategies to interrupt the feedback loops are essential for mitigating these issues.
5.2. Unified Mitigation Framework
Current methods primarily address individual instances of bias or unfairness, but in the future, there is a need to consider unified solutions. This is because various types of bias and unfairness are not isolated; they are interconnected, and presenting a unified framework can facilitate a deeper understanding of these relationships, enabling methods for addressing different types of bias and unfairness to complement each other. Our proposed unified perspective in this survey offers a potential direction to address these issues simultaneously.
5.3. Theoretical Analysis and Guarantees
The current exploration of bias and unfairness within the intersection between LLMs and IR systems has predominantly been through empirical studies. However, there is a critical need for robust theoretical analysis to augment these empirical findings. The definition and mitigation of bias and unfairness from a unified distribution matching perspective in this survey offer promising groundwork for further theoretical exploration. Future efforts should focus on developing more rigorous analytical frameworks that not only explore bias and unfairness phenomena but also provide theoretical insights for the development of effective mitigation strategies.
5.4. Better Benchmarks and Evaluation
Most benchmarks currently utilized to study bias and unfairness at the intersection of LLMs and IR systems are designed within simulated environments. There is a crucial need for collecting large-scale, real-world datasets to enhance the evaluations and broaden research horizons. Additionally, as LLMs increasingly draw upon existing online data to train subsequent generations, dynamic benchmarks are needed to be constructed. Furthermore, current research often leverages distinct methods for evaluating different types of bias and unfairness, which hinder the development of unified mitigation models. Consequently, future work can focus on exploring a systematic evaluation protocol for different bias and unfairness issues.
6. Conclusion

### v2

@@ 3.2.2. Popularity Bias
$\bullet$ Definition. LLM-based IR models tend to prioritize candidate documents or items with high popularity levels.
@@ 3.2.3. Instruction-Hallucination Bias
$\bullet$ Definition. Content generated by LLM-based IR models may deviate from the instructions provided by users.
@@ 3.2.4. Context-Hallucination Bias
$\bullet$ Definition. LLMs-based IR models may generate content that is inconsistent with the context.
There have been many studies showing that LLMs run the risk of generating content that is inconsistent with the context, especially in scenarios where the context is very long and multi-turn responses are needed (Liu et al., 2024; Qiu et al., 2024; Bai et al., 2024; Li et al., 2023f). When using LLMs as recommenders in scenarios with complex context such as multiple rounds of conversation history and user portrait information, LLMs may give the items that are contradictory with the conversation history and user preferences. This emerges when LLMs cannot understand the context or fail to maintain consistency throughout the conversation (Zhang et al., 2023d), which is mainly because LLMs still have limitations in processing long texts (Liu et al., 2024). Therefore, the main research to mitigate the context-hallucination bias focuses on improving the memory and processing capabilities of LLMs for long texts (Xu et al., 2023c; Chen et al., 2023c; Wang et al., 2024).
@@ 3.3. Bias in Result Evaluation
3.3.1. Selection Bias
A primary challenge in utilizing LLMs as evaluators is that they are sensitive to the order/ID tokens of candidate responses, a phenomenon known as selection bias:
$\bullet$ Definition. LLM-based evaluators may favor the responses at specific positions or with specific ID tokens.
For example, Zheng et al. (2023c) have demonstrated that gpt-3.5-turbo exhibits a preference for choice “C”, while llama-30B shows a preference for choice “A” across various benchmarks. Other works have revealed a common tendency among LLMs to favor responses positioned at the first position  (Wang et al., 2023b; Pezeshkpour and Hruschka, 2023; Koo et al., 2023; Zheng et al., 2024; Chen et al., 2024a). Selection bias may stem from an imbalance in how answers of different positions or with distinct ID options are represented in the training data (Zheng et al., 2024, 2023c). Further investigation reveals that this bias tends to amplify when LLM-based evaluators are uncertain about the prediction between the top-ranked choices(Wang et al., 2023b; Zheng et al., 2023c).
To address selection bias, several approaches have been explored: (1) Prompting: Simple prompt-based methods, such as incorporating few-shot examples or employing chain-of-thought and reference-guided judgment, have been proposed (Zheng et al., 2024; Pezeshkpour and Hruschka, 2023; Kim et al., 2023). (2) Data Augmentation: Strategies like position or token switching aim to eliminate selection bias by diversifying the evaluation context (Zheng et al., 2024; Chern et al., 2024; Zheng et al., 2023b; Wang et al., 2023b; Pezeshkpour and Hruschka, 2023). While these methods can enhance the robustness of evaluations, they are often time-consuming and costly. (3) Rebalancing: Zheng et al. (2023c) utilized probability decomposition techniques to estimate the prior distribution of specific positions or tokens associated with responses, which helps in aligning evaluations closer to objective standards. Wang et al. (2023b) have developed a calibration framework that integrates Human-in-the-Loop to calculate balanced position diversity entropy for final selection. Despite the availability of these strategies, effectively mitigating selection bias in LLM-based evaluators remains a challenge, requiring further research for more efficient solutions.
@@ 3.3.2. Style Bias
$\bullet$ Definition. LLM-based evaluators may favor the responses with specific styles (e.g., longer responses).
For instance, several studies have identified a clear preference in LLMs for longer responses over shorter ones, emphasizing form over the actual quality of the content (Wu and Aji, 2023; Zheng et al., 2024; Liu et al., 2023b; Saito et al., 2023). Moreover, Chen et al. (2024a) have observed an inclination towards content with visually engaging elements, such as emojis or references, even when such content may include factual errors. Huang et al. (2024) suggest that this bias may stem from the training process of LLMs, which emphasizes generating fluent and verbose responses, thereby inadvertently leading them to prefer these characteristics when employed as evaluators.
@@ 3.3.3. Egocentric Bias
With LLMs being extensively utilized in the development of IR models, egocentric bias has emerged as a new bias during the automated evaluation conducted by LLMs (Liu et al., 2023b, a; Koo et al., 2023; Xu et al., 2024e), which can be defined as follows:
$\bullet$ Definition. LLM-based evaluators prefer the responses generated by themselves or LLMs from the same family.
@@
Data Collection User Unfairness (Ghanbarzadeh et al., 2023; Zhang et al., 2023a; Xu et al., 2023d; Lu et al., 2020; Sun et al., 2022b; Ung et al., 2021) (Raffel et al., 2020; Ngo et al., 2021) (Orgad and Belinkov, 2022; Deldjoo and di Noia, 2024) (Qian et al., 2019; Bordia and Bowman, 2019; Huang et al., 2019) (Fang et al., 2024)
@@
Item Unfairness (Grosse et al., 2023) (Shi et al., 2023; Akyürek et al., 2022) (Wang et al., 2023h; Zhang et al., 2023b, e; Wang et al., 2023c; Sander et al., 2024)
@@ 4. Cause and Mitigation of Unfairness
4.1. Fairness Concepts
Sociological researches acknowledge multiple cultural variations in perceptions of fairness (Tyler and Allan Lind, 2002; Tyler and Smith, 1995). In IR systems, achieving fairness often entails ensuring that the retrieved documents or recommended items align with cultural values (Mehrabi et al., 2021; Li et al., 2023a), including principles such as gender equality (Fang et al., 2024; Zhang et al., 2023a; Xu et al., 2023d; Li et al., 2023a), addressing disadvantages (Xu et al., 2023a; Patro et al., 2020), and avoiding discriminatory language (Wen et al., 2023; Gallegos et al., 2023).
Researchers have revealed that various multi-stakeholders involved in IR systems (Abdollahpouri et al., 2020; Chen et al., 2023a), such as users and items, often have distinct perspectives on fairness considerations. In IR, user fairness and item fairness are often associated with two sociological concepts: equality and distributive justice (Xu et al., 2023a).
$\bullet$ User Fairness. Everyone should be treated the same and provided the same resources to succeed. This implies that the IR systems should deliver equitable and non-discriminatory information services to different users.
$\bullet$ Item Fairness. The resources should be equally distributed based on needs. This implies that the IR systems should afford more opportunities (e.g., exposures) to weaker items, striving to equalize the opportunities across diverse items.
@@ 4.2.1. User Unfairness
In the context of user unfairness, one primary cause stems from the existing taxonomic, discriminatory, and offensive content in the training data, disproportionately affecting specific groups. The inclusion of these contents within the existing material can be attributed to historical and cultural reasons (Ntoutsi et al., 2020; Zhuo et al., 2023; Beukeboom et al., 2019), or they may be generated by LLMs (Fang et al., 2024). For example, when training LLMs, it is common to encounter discriminatory content (Abid et al., 2021; Fang et al., 2024; Das et al., 2024). Discrimination against certain groups can also stem from unbalanced data collection, where the insufficient representation of diverse perspectives leads to unfair outcomes (Ghanbarzadeh et al., 2023; Mehrabi et al., 2021). The presence of unbalanced data can contribute to the perpetuation of historical and cultural stereotypes (Ellemers, 2018; Haines et al., 2016; Ladhak et al., 2023) or systematic influences (Chen et al., 2023a).
Previous works have employed various methods to mitigate unfairness during data collection by redistributing the existing document corpus. Specifically, some work (Ghanbarzadeh et al., 2023; Zhang et al., 2023a; Xu et al., 2023d; Lu et al., 2020) create matched pairs (e.g., male or female) to ensure a more equitable dataset and other methods (Dixon et al., 2018; Sun et al., 2022b; Ung et al., 2021) add non-toxic examples for groups. Other approaches (Orgad and Belinkov, 2022; Deldjoo and di Noia, 2024) suggest using down-weighting samples containing social group or discriminated information. Moreover, other studies (Raffel et al., 2020; Ngo et al., 2021) propose to filter out discriminated or taxonomic content from web-scale datasets. Finally, instruction fine-tuning or RLHF has also been shown to be effective in promoting fairness (Touvron et al., 2023; Lu et al., 2020).
@@ 4.2.2. Item Unfairness
For item unfairness, one primary cause is likely unbalanced data collection, where the insufficient representation of certain items leads to disparities in the IR process (Jiang et al., 2024a). Another reason raised is that LLMs cannot only retrieve existing items but also generate new items, contributing to the potential introduction of novel content and perspectives (Das et al., 2024; Laverghetta Jr and Licato, 2023; Guenole et al., 2024; Götz et al., 2023; Zu et al., 2023). However, these newly generated items may still encounter discrimination issues (Laverghetta Jr and Licato, 2023; Guenole et al., 2024).
@@ 4.3.1. User Unfairness
When adapting LLMs as information retrievers, researchers have observed that the extensive knowledge gained during pre-training may introduce risks of user unfairness (Xu et al., 2023d; Zhang et al., 2023a; Ghanbarzadeh et al., 2023; Deldjoo and di Noia, 2024), highlighting the need for careful consideration and mitigation strategies in deploying such models. Studies (Zhang et al., 2023a; Deldjoo and di Noia, 2024; Ghanbarzadeh et al., 2023) have shown that utilizing explicit user-sensitive attributes like gender or race in LLMs may lead to the generation of discriminated recommendation results or unfair answers to specific questions. Moreover, it has been observed that LLMs can learn implicit attributes, such as user names and email addresses, and utilize them to generate discriminated content (Xu et al., 2023d; Wolfe and Caliskan, 2021).
To address user unfairness in the model development phase, prior research has suggested mitigating unfairness during the fine-tuning process. Some studies (Deldjoo and di Noia, 2024) investigate how different intersectional prompts affect recommendation fairness and UP5 (Hua et al., 2023; Yang et al., 2023b) propose to conduct prompt tuning to get an effective fair-aware prompt. Han et al. (2021); Zayed et al. (2023) proposes to set different weights of loss for different samples containing discriminated content. Meanwhile, various studies (Liu et al., 2019; Woo et al., 2023; Park et al., 2023; Wang et al., 2023a; Garimella et al., 2021; Zhou et al., 2023) explore the incorporation of a fairness-aware regularizer to assist LLMs-based models in generating more equitable content. Wang et al. (2021) proposes to remove unfair information from LLMs-based embeddings by generating adversarial examples. There are also some works (Wang et al., 2021; Meade et al., 2023; Schramowski et al., 2022; Shuster et al., 2022) that recommend employing a filtering-list approach or comparing model outputs with safe samples to proactively prevent the inclusion of discriminatory words and enhance the fairness of the generated content.
@@ 4.4.2. Item Unfairness
In evaluating item unfairness, when turning from discriminant style to generation style, a primary concern arises from attributing credit to the generated items (Shi et al., 2023; Grosse et al., 2023; Akyürek et al., 2022; Wang et al., 2023c; Sander et al., 2024), as achieving item fairness necessitates tracing this credit back to the item provider for a comprehensive assessment.
@@
To tackle these challenges, several studies (Wang et al., 2023h; Zhang et al., 2023b, e) have developed IR agents that use world knowledge to fairly distribute item credits, aiming to reduce biases and enhance fairness. Shi et al. (2023) employs the MIN-KProb technique to check whether an item exists in the training data. Meanwhile, Grosse et al. (2023); Akyürek et al. (2022) leverage influence functions and embedding similarities for item credit assessment. Additionally, other research (Wang et al., 2023c; Sander et al., 2024) explores watermarking techniques for item tracking.
5. Conclusion and Future Directions